# Optimizing VarLiNGAM for Scalable and Efficient Time Series Causal Discovery

### v1

@@ 6.2.1 Using GPUs on current alogrithm
Considering the advantages of GPUs and CPUs, hybrid architectures can be used. For example, the GPU can be used to handle batch computations on large-scale data sets while letting the CPU handle lighter control logic and non-parallel tasks. This maximises both benefits and improves overall performance. When using GPUs, we could investigate specific ways to optimise the speed of algorithms using CUDA. During causal graph learning, targeted CUDA optimisation of intensive computational tasks is required to reduce the time of each computational step.
@@ 6.2.2 Optimising pruning step in VarLiNGAM
The Pruning process is a very crucial part of VarLiNGAM, aiming to simplify the causal graph by removing irrelevant or redundant edges. However, existing pruning methods may still have space for optimisation in terms of efficiency and accuracy.

### v2

@@ 6.2.1 Using GPUs on current alogrithm
Considering the advantages of GPUs and CPUs, hybrid architectures can be used. For example, the GPU can be used to handle batch computations on large-scale data sets while letting the CPU handle lighter control logic and non-parallel tasks. This maximizes both benefits and improves overall performance. When using GPUs, we could investigate specific ways to optimise the speed of algorithms using CUDA. During causal graph learning, targeted CUDA optimization of intensive computational tasks is required to reduce the time of each computational step.
@@ 6.2.2 Optimising pruning step in VarLiNGAM
The Pruning process is a very crucial part of VarLiNGAM, aiming to simplify the causal graph by removing irrelevant or redundant edges. However, existing pruning methods may still have space for optimization in terms of efficiency and accuracy.